OpenStudy (anonymous):

Addison is mixing hummingbird food to fill her feeders. Ideally, the hummingbird food should contain 25% sugar. Addison remembers that she has 8 ounces of leftover hummingbird food that contains 35% sugar. She decides to mix some diluted food to add to the leftover food. If the diluted food contains 10% sugar, how much should she add to the food with 35% sugar to get the ideal hummingbird food with 25% sugar?

OpenStudy (anonymous):

x = oz of 10% y = oz of 35% x + y = 8 10x + 35y = 25*8

OpenStudy (anonymous):

how would you get the answer?

OpenStudy (anonymous):

Solve the system

OpenStudy (anonymous):

i don't understand how to

OpenStudy (anonymous):

there are three different methods, you should have learned at least one of them

OpenStudy (anonymous):

i haven't

OpenStudy (anonymous):

i need help really bad @rachelfreak123

OpenStudy (anonymous):

hello??? @rachelfreak123